# Towards Black-Box Membership Inference Attack for Diffusion Models

### v4

@@ 1 Introduction
Recently, there has been a surge in the popularity of generative models, with diffusion models in particular, gaining huge attention within the AI community (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Song et al., 2020b). These models have demonstrated remarkable capabilities across various tasks, including unconditional image generation (Ho et al., 2020; Song et al., 2020a), text-to-image generation (Rombach et al., 2022; Yu et al., 2022; Nichol et al., 2021) and image-to-image generation (Saharia et al., 2022a). This surge has given rise to powerful AI art models such as DALL-E 2 (Ramesh et al., 2022), Stable Diffusion (Rombach et al., 2022), and Imagen (Saharia et al., 2022b). AI-generated art holds a promising future and is expected to have a widespread impact.
@@
A series of studies have been conducted on membership inference attacks against diffusion models. Hu & Pang (2023) was the first to examine this issue, utilizing the loss function values of diffusion models to determine whether an image is in the training set. Duan et al. (2023) and Kong et al. (2023) extended this work by relaxing the assumptions about model access requirements.
@@ 4.2 Algorithm
[…]
@@
Denote $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, we assume that the denoise model is expressive enough(the neural network’s dimensionality significantly exceeds that of the image data) such that for the input $x_{0}\in\mathbb{R}^{d}$ and time step $t\in\mathcal{T}$, the Jacobian matrix $\nabla_{\theta}\epsilon_{\theta}(x_{t},t)$ is full rank $(\geq d)$. This suggest that the model can adjust the predicted noise $\epsilon_{\theta}(x_{t},t)$ locally in any direction. Then for a well trained model, we would have $\nabla_{\theta}L(\theta)=0$. This implies
@@
This is intuitive because if the noise prediction from the neural network exhibits high bias, the network can adjust to fit the bias term, further reducing the training loss.
@@
Our algorithm classifies a sample as being in the training set if $D(x,\hat{x})$ is smaller than a threshold $\tau$, where $D(x,\hat{x})$ represents the difference between the two images. It can be calculated using traditional functions, such as the SSIM metric (Wang et al., 2004). Alternatively, we can train a neural network as a proxy. In Section 5, we will introduce the details of $D(x,\hat{x})$ used in our experiment.
@@
  return ”YES” if the distance between the two images $D(x,\hat{x})$ is less than $\tau$, otherwise ”NO”.
@@
The theorem suggests that averaging the randomly variational data from a training set will result in a smaller relative error with high probability when we use a large $n$. We defer the proof of this theorem to Appendix C.
@@ 5 Experiments
In this section, we evaluate the performance of our methods across various datasets and settings. We follow the same experiment setup in previous papers (Duan et al., 2023; Kong et al., 2023). The detailed experimental settings, including datasets, models, and hyper-parameter settings can be found in Appendix A.
@@ 5.2 MIA with DDIM Models
For the difference function $D(x,\hat{x})$, following the setup in (Duan et al., 2023), we take the pixel-wise absolute value of $x-\hat{x}$ to obtain a difference vector $v$ for each image. Using the ResNet-18 network (He et al., 2016) and denoting it as $f_{R}$, we perform binary classification on these difference vectors. We use 20% of the data as the training set and obtained the label of each difference vector being classified as a member or nonmember. The difference function is obtained by the negated value of the probability outputed by the neural network predicting as member: $D(x,\hat{x})=-f_{R}(v)$.
@@ 5.3 MIA with Diffusion Transformers
We train a diffusion transformer model on the ImageNet (Deng et al., 2009) dataset following the setup of (Peebles & Xie, 2023). We randomly select 100,000 images from the ImageNet training set to train the model with resolutions of either $128\times 128$ or $256\times 256$. For the membership inference attack setup, 1000 images are randomly chosen from our training set as the member set, and another 1000 images are randomly selected from the ImageNet validation set as the non-member set. We fix the diffusion step at $t=150$ and the DDIM step at $k=50$, and independently call the variation API 10 times to take the average of the output images as $\hat{x}$.
We use (Matsumoto et al., 2023), (Duan et al., 2023), (Kong et al., 2023) as our baseline methods. Since these work did not study the case of Diffusion Transformers, we integrate their algorithms into the DiT framework for evaluation. For the difference function $D(x,\hat{x})$, following the setup in (Duan et al., 2023; Kong et al., 2023), we take the L2 norm of $x-\hat{x}$ to measure the differences between two image.The results, presented in Table 2, demonstrate that our method outperforms baseline algorithms and does not require access to the Vision Transformer.
@@ 5.4 MIA with the Stable Diffusion Model
For the difference function $D(x,\hat{x})$, since the images in these datasets better correlate with human visual perception, we directly use the SSIM metric (Wang et al., 2004) to measure the differences between two images. The results, presented in Table 3, demonstrate that our methods achieve high accuracy in this setup, outperforming baseline algorithms by approximately 10%. Notably, our methods do not require access to U-Net.
@@ Appendix A Datasets, Models and Hyperparameters
For the Diffusion Transformers, we train the model on the ImageNet (Deng et al., 2009) dataset, using 1000 image generation steps ($T=1000$), while for other training hyperparameters, we follow the setup of (Peebles & Xie, 2023). We randomly select 100,000 images from the ImageNet training set to train the model at resolutions of either $128\times 128$ or $256\times 256$. For the $128\times 128$ image size, we use 160,000 training iterations. For the $256\times 256$ image size, we use 300,000 training iterations. These numbers of training iterations are chosen to ensure the generation of high-quality images.
For the membership inference attack setup, 1000 images are randomly selected from our training set as the member set, and another 1000 images are randomly chosen from the ImageNet validation set as the non-member set. We fix the diffusion step at $t=150$ and the DDIM step at $k=50$, and independently call the variation API 10 times to average the output images as $\hat{x}$.
@@ Appendix C Proof
In this section, we implement the proof of the theorem. See 4.2
@@
We denote $x^{i}$ as the $i$-th output image of $x$ using the variation API. We denote the $i$-th Gaussian noise as $\epsilon^{i}$ and the forward process incur $x^{i}_{t}=\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i}$.
@@
If we denote the random variable $X^{i}=(X^{i}_{1},X^{i}_{2},\ldots,X^{i}_{d})$ represents $\epsilon^{i}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},t)$, then we consider $S^{i}_{n}=\sum_{j=1}^{n}X^{i}_{j}$. From the assumption we know that all the $X^{i}_{j}$ are random variables with zero expectation and finite cumulant-generating function $\Psi_{X}(s)=\log E[e^{sX}]<+\infty$. Using the theorem of Chernoff-Cramer method for sums of IID RV (Durret, 2010), we get the following probability inequality for any $\beta>0$:
@@
where $\beta>0$ and $\Psi^{*}_{X}(y)=\sup_{s>0}(sy-\Psi_{X}(s))$ is the Fenchel-Legendre dual of the cumulant-generating function $\Psi_{X}$.
@@
So averaging the randomly reconstructed data from a training set will result in a smaller reconstruction error with high probability of $\Theta(1-\exp(-n))$ when we use a large $n$.

### v5

@@ 1 Introduction
Recently, there has been a surge in the popularity of generative models, with diffusion models in particular, gaining huge attention within the AI community (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Song et al., 2020b). These models have demonstrated remarkable capabilities across various tasks, including unconditional image generation (Ho et al., 2020; Song et al., 2020a), text-to-image generation (Rombach et al., 2022; Yu et al., 2022; Nichol et al., 2021) and image-to-image generation (Saharia et al., 2022a). This surge has given rise to powerful AI art models such as DALL-E 2 (Ramesh et al., 2022), Stable Diffusion (Rombach et al., 2022), and Imagen (Saharia et al., 2022b). AI-generated art has a promising future and is expected to have widespread impact.
@@
A series of studies have been conducted on membership inference attacks against diffusion models. Hu & Pang (2023) was the first to examine this issue, utilizing the loss function values of diffusion models to determine whether an image is in the training set. Duan et al. (2023) and Kong et al. (2023) extended this work by relaxing assumptions about the access requirements of the model.
@@ 4.2 Algorithm
In this section, we present the intuition of our algorithm. We denote $\|\cdot\|$ as the $L_{2}$ operator norm of a vector and $\mathcal{T}=\{1,2,\ldots,T\}$ as the set of diffusion steps. The key insight is derived from the training loss function of a fixed sample $x_{0}$ and a time step $t\in\mathcal{T}$:
@@
Denote $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, we assume that the denoise model is expressive enough (the neural network’s dimensionality significantly exceeds that of the image data) such that for the input $x_{0}\in\mathbb{R}^{d}$ and time step $t\in\mathcal{T}$, the Jacobian matrix $\nabla_{\theta}\epsilon_{\theta}(x_{t},t)$ is full rank $(\geq d)$. This suggests that the model can adjust the predicted noise $\epsilon_{\theta}(x_{t},t)$ locally in any direction. Then for a well trained model, we would have $\nabla_{\theta}L(\theta)=0$. This implies
@@
This is intuitive because if the neural network noise prediction exhibits a high bias, the network can adjust to fit the bias term, further reducing the training loss.
@@
Our algorithm classifies a sample as being in the training set if D(x,x^)D($\hat{x}$) is smaller than a threshold $\tau$, where D(x,x^)D($\hat{x}$) represents the difference between the two images. It can be calculated using traditional functions, such as the SSIM metric (Wang et al., 2004). Alternatively, we can train a neural network as a proxy. In Section 5, we will introduce the details of D(x,x^)D($\hat{x}$) used in our experiment.
@@
return ”YES” if the distance between the two images D(x,x^)D($\hat{x}$) is less than $\tau$, otherwise ”NO”.
@@
The theorem suggests that averaging the random variational data from a training set will result in a lower relative error with high probability when we use a large $n$. We defer the proof of this theorem to Appendix C.
@@ 5 Experiments
In this section, we evaluate the performance of our methods across various datasets and settings. We follow the same experiment setup in previous papers (Duan et al., 2023; Kong et al., 2023). The detailed experimental settings, including datasets, models, and hyper-parameter settings can be found in Appendix A [^1] Code is available at https://github.com/lijingwei0502/diffusion_mia..
@@ 5.2 MIA with DDIM Models
For the difference function D(x,x^)D($\hat{x}$), following the setup in (Duan et al., 2023), we take the pixel-wise absolute value of $x-\hat{x}$ to obtain a difference vector $v$ for each image. Using the ResNet-18 network (He et al., 2016) and denoting it as $f_{R}$, we perform binary classification on these difference vectors. We use 20% of the data as the training set and obtained the label of each difference vector being classified as a member or nonmember. The difference function is obtained by the negated value of the probability outputed by the neural network predicting as member: $D(x,\hat{x})=-f_{R}(v)$.
@@ 5.3 MIA with Diffusion Transformers
We train a diffusion transformer model on the ImageNet (Deng et al., 2009) dataset following the setup of (Peebles & Xie, 2023). We randomly select 100,000 images from the ImageNet training set to train the model with resolutions of either $128\times 128$ or $256\times 256$. For the membership inference attack setup, 1000 images are randomly chosen from our training set as the member set, and another 1000 images are randomly selected from the ImageNet validation set as the non-member set. We fix the diffusion step at $t=150$ and the DDIM step at $k=50$, and we independently call the variation API 10 times to take the average of the output images as $\hat{x}$.
We use (Matsumoto et al., 2023), (Duan et al., 2023), (Kong et al., 2023) as our baseline methods. Since these work did not study the case of Diffusion Transformers, we integrate their algorithms into the DiT framework for evaluation. For the difference function D(x,x^)D($\hat{x}$), following the setup in (Duan et al., 2023; Kong et al., 2023), we take the L2 norm of $x-\hat{x}$ to measure the differences between two image.The results, presented in Table 2, demonstrate that our method outperforms baseline algorithms and does not require access to the Vision Transformer.
@@ 5.4 MIA with the Stable Diffusion Model
For the difference function D(x,x^)D($\hat{x}$), since the images in these datasets better correlate with human visual perception, we directly use the SSIM metric (Wang et al., 2004) to measure the differences between two images. The results, presented in Table 3, demonstrate that our methods achieve high accuracy in this setup, outperforming baseline algorithms by approximately 10%. Notably, our methods do not require access to U-Net.
@@ Appendix A Datasets, Models and Hyperparameters
For the Diffusion Transformers, we train the model on the ImageNet (Deng et al., 2009) dataset, using 1000 image generation steps ($T=1000$), while for other training hyperparameters, we follow the setup of (Peebles & Xie, 2023). We randomly select 100,000 images from the ImageNet training set to train the model at resolutions of either $128\times 128$ or $256\times 256$. For the $128\times 128$ image size, we use 200,000 training iterations. For the $256\times 256$ image size, we use 300,000 training iterations. These numbers of training iterations are chosen to ensure the generation of high-quality images.
For the membership inference attack setup, 1000 images are randomly selected from our training set as the member set, and another 1000 images are randomly chosen from the ImageNet validation set as the non-member set. We fix the diffusion step at $t=100$ and the DDIM step at $k=50$, and independently call the variation API 10 times to average the output images as $\hat{x}$.
@@ Appendix C Proof
In this section, we present the proof of the theorem. See 4.2
@@
We denote $x^{i}$ as the $i$-th output image of $x$ using the variation API. We denote the $i$-th Gaussian noise as $\epsilon^{i}$ and the forward process yields $x^{i}_{t}=\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i}$.
@@
If we denote the random variable $X^{i}=(X^{i}_{1},X^{i}_{2},\ldots,X^{i}_{d})$ represents $\epsilon^{i}-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon^{i},t)$, then we consider $S^{i}_{n}=\sum_{j=1}^{n}X^{i}_{j}$. From the assumption we know that all the $X^{i}_{j}$ are i.i.d. random variables(with the same distribution of $X_{i}$ in Theorem 4.2) with zero expectation and finite cumulant-generating function $\Psi_{X_{i}}(s)=\log E[e^{sX_{i}}]<+\infty$. Using the theorem of Cramér–Chernoff method for sums of i.i.d. random variables (Durret, 2010), we get the following probability inequality for any $\beta>0$:
@@
where $\beta>0$ and $\Psi^{*}_{X_{i}}(y)=\sup_{s>0}(sy-\Psi_{X_{i}}(s))$ is the Fenchel-Legendre dual of the cumulant-generating function $\Psi_{X_{i}}$.
@@
So averaging the randomly reconstructed data from a training set will result in a smaller reconstruction error with high probability of $1-\mathcal{O}(\exp(-n))$ when we use a large $n$.